# GlucOS: Security, correctness, and simplicity for automated insulin delivery

### v3

@@ 3.2. Model principles
Our contribution is not in the specific algorithms we use, but rather in the insight around insulin delivery being equal for all correct algorithms and how we use this insight for security. In this section, we outline the algorithms we use in our current implementation and define the principles for how to apply security to any predictive insulin dosing algorithms.
Algorithmically, we use a PID controller (franklin2002feedback, ) from feedback control for our reactive safe model, and we use a deep neural-network from the literature for meal predictions for our predictive ML model (mosquera2023enabling, ). Although we use this model in our current implementation, our system supports any predictive model. PID controllers are used in automated insulin delivery systems (weinzimer2008fully, ; kushner2018data, ) and research predictive ML models for insulin delivery have been well covered in the literature (Zhu2022, ; Jacobs, ; Zhang2021, ; Josep2020, ; RL, ; Tyler2022, ).
The design of the reactive safe model is crucial for algorithmic security, as it serves as the anchor for the security mechanism. The ideal reactive safe model should be safe, easy to understand, and have a formally verified implementation (Section 7).
@@
To ensure the safety of this model, we base its calculations solely on facts observed automatically from the CGM and the insulin pump. Given these high-quality sources of data, and a simple insulin absorption physiological model, we can calculate how much insulin an individual needs and how much insulin they have injected but have not yet activated. From this calculation, we know how much insulin to inject or withhold.
@@
These calculations are standard calculations that people who inject insulin manually use to determine dosing, making them easy for people who manage T1D or their medical care team to understand. The difference in our system is that we run these calculations automatically and every five minutes to adjust to the latest sensor readings. This basic formulation is a standard formulation used in most automated insulin delivery systems.
In contrast, predictive models have no restrictions since their outputs are constrained by the reactive safe model and our security mechanism. The ideal predictive model anticipates future metabolic states and adjusts insulin proactively to avoid glycemic imbalance.
In our evaluation, our predictive model outperforms the reactive safe model both in simulation (Section 9) and for the individual that uses our system for their real-world insulin dosing (Section 10), showing why it is important for automated insulin delivery systems to have a practical way to adopt ML.
4. Driver security: Withstanding malicious pump drivers
Given this complexity, there is a strong possibility for bugs or vulnerabilities to be present in existing pump drivers. These vulnerabilities can harm or even lead to the death of individuals since pump drivers control insulin pumps. For example, consider a scenario where the user decides to inject 3 units (3U) of insulin using the BioKernel. The BioKernel would relay this to the pump through abstractions provided by the pump driver. If the pump driver inadvertently or intentionally issues this command twice, 6U would be delivered, which could endanger the user by dosing too much insulin.
Our security goals for GlucOS’s driver security are to protect users from inadvertent bugs in pump drivers and safeguard against potentially malicious pump drivers. To achieve these goals, we introduce a shim between the pump driver and the pump, which validates all communications between them. We design our driver security mechanism to interpose on existing pump driver-pump communication instead of directly programming the pump. This design ensures that our trusted computing base remains small and simple by not including the complexity of pump driver implementations, thereby enabling formal verification. This design also provides flexibility with altering or introducing new functionality to pump drivers.
The most interesting aspect of our driver security mechanism was having to overrule our principle of not programming the pump in certain cases in the interest of protecting the individual. These cases involve commands pertaining to canceling ongoing insulin delivery. We could not directly adopt the design of other device driver security mechanisms (ref_validation, ; tang2010trust, ; expressos, ) for our driver security mechanism since these cases are specific to insulin pumps.
We focus our security efforts on pump drivers because they are complex (14.5k LoC), multi-threaded, event-driven, and use two-way communication between the pump and the driver. In contrast, CGM drivers are simple (1.9k LoC) and only read sensor values from the device. In our future work, we will formally verify our CGM driver using the same techniques we use for the insulin delivery path (Section 7).
4.1. Driver security mechanism design
Figure 4 provides an overview of our driver security mechanism. The BioKernel registers high-level commands, like “deliver 2U of insulin” with our driver security mechanism, and sends them to the pump driver. The pump driver converts these commands into I/O messages to program the pump. The driver security mechanism then checks to make sure that these I/O messages are consistent with the previously registered high-level pump commands before forwarding them to the pump hardware. We only validate commands for insulin delivery since they pose the biggest threats to the user. On detecting added, modified or dropped commands from the pump driver, the driver security mechanism rejects further commands and transitions the system to a manual mode where the user takes control.
Dropped commands pertaining to canceling ongoing insulin delivery requires separate consideration since excess insulin is dangerous and cannot be withdrawn once it has been delivered. Users will cancel insulin delivery if they accidentally issued an insulin delivery command or incorrectly set the insulin dosage. In cases where a malicious driver drops a cancel command, the driver security mechanism directly issues I/O messages to the pump to cancel insulin delivery before transitioning the pump to a manual mode. We make the tradeoff of increasing the size of our trusted computing base to handle these cases in the interest of protecting the user. Fortunately, there are only two commands to cancel insulin delivery in our current implementation, so the increase to our trusted computing base is small.
5. End-to-end security: Biological invariant
When implementing security systems for humans, even the perfect algorithmic and driver security mechanisms are unable to handle all situations due to changes in human physiology. A key novelty in our overall approach is our introduction of an end-to-end check that ensures the BioKernel’s glucose calculations against observations of the user’s metabolic state. If there is a large mismatch between our calculations and what we observe, then it means that our calculations are wrong. These deviations can happen from drastic changes in the user’s glucose absorption, like during exercise when the body uses glucose without needing insulin (borghouts2000exercise, ), or if the pump insertion site hits a vein and delivers insulin directly into the blood stream (steyn2023accelerated, ). Regardless of the cause, the end-to-end check detects these conditions and prevents harm to the user. We call this safety check that accounts for changing human physiology a biological invariant.
In our implementation, we check that changes in glucose levels that we observe from CGM readings and the individual’s insulin delivery history are consistent with our theoretical calculations of how much the person’s glucose should change given these same inputs. If their glucose is dropping by more than our theoretical calculations show, then they are more sensitive to insulin then our system believes they should be. We set a threshold of 30 mg/dl per hour, for the maximum allowable divergence based on clinical evaluations showing detection of significant glycemic deviations above this level (30mg_1, ; 30mg_2, ).
We take a three-tiered approach to dealing with deviations in our glucose change calculations. First, for small differences of less than 30 mg/dl, we accumulate errors and adjust the target glucose level to compensate, clamping the adjustment to 10 mg/dl. For example, if the individual’s insulin sensitivity is elevated and their target glucose level is 90 mg/dl, GlucOS will update their target to 100 mg/dl dynamically to adjust. Second, when the differences are 30 mg/dl or greater, we shut off insulin delivery. Our rationale is that the individual is more sensitive to insulin than our models believe they should be, thus we want to avoid overdosing. If their difference goes back to within the expected range, normal execution resumes. Third, if they remain at 30 mg/dl or more for more than two hours, we alert the individual and transition the system into a manual mode that requires the user to dose insulin manually.
@@ 6. Last line of defense: Humans
In this section, we describe our design and implementation for incorporating humans as a critical component of our multi-layered security system. While our algorithmic, system-level, and biological invariant defenses form the primary barriers against attacks, humans serve as the last line of defense, capable of mitigating attacks that manifest as glycemic imbalance. We discuss our most interesting findings around preferences and the architectural iterations we make to our design based on our real-world experiences.
@@ 6.1. Predictive alerts for humans
When an individual’s glucose levels are not within the range of healthy values, they need to intervene to bring them back in range. Low glucose, or hypoglycemia, occurs from too much insulin, and there is no way to remove insulin that the automated insulin delivery system has already delivered. To recover from hypoglycemia, people need to eat sugar. Given that people with T1D are already active participants in managing glycemic imbalance, we focus on how we can use this fact to our advantage to create a more comprehensive security system.
@@
We use predictive yet deterministic mobile notifications to alert people when they are likely to experience hypoglycemia before they reach it (Figure 5), as hypoglycemia presents the most immediate health risk and could be an indicator of an attack. Alerts for hypoglycemia are not new. All CGM software comes with non-predictive hypoglycemia alerts by default (dexcom_alerts, ) and glucose prediction algorithms have been well covered in the literature (Josep2020, ; Zhang2021, ; arora2023multivariate, ). Our contribution is using predictive alerts before someone reaches hypoglycemia as a security defense.
In a user study (Section 10.4), we reduced the amount of time six people spent in hypoglycemia as a result of using GlucOS, despite them already having CGM alerts set up prior to our study. These results demonstrate a reduction in hypoglycemia for the non-attack case, but the most dangerous successful attacks will result in hypoglycemia. Thus, these results suggest that predictive alerting can be an effective security measure.
Using a simple and deterministic algorithm is important since it runs as a part of our trusted computing base. Algorithmically, we predict hypoglycemia by doing a least squares fit for the last 20 minutes of CGM readings and extrapolate forward 15 minutes. We trigger alerts if our prediction shows hypoglycemia is likely in the next 15 minutes.
The most interesting finding from our study was that personalization was important. We provided people with the ability to set the prediction threshold and specify how often they wanted the alert to repeat while they remained out-of-range. Five of the six participants customized their alerts and the participant noted that the default values we picked coincided with their alerting preferences. In our first implementation we did not provide the ability to customize their security settings and we tried to make the usability vs security tradeoff algorithmically and automatically. But in the end, simplifying the system and allowing users to personalize their security policy was important for our participants.
@@ 6.2. Architectural implications
Moving the alerting logic into the BioKernel did have security implications. It adds complexity to our trusted computing base, but we carefully manage this tradeoff. We minimize the added complexity by using traditional, well-understood algorithms for prediction rather than more complex ML models and reuse existing functionality within the BioKernel when possible to keep the implementation simple. This approach allowed us to maintain a high level of confidence in the security and correctness of our alerting system while still providing the necessary functionality.
@@ 7. Formal verification
To provide additional assurance that our implementation is correct, we port our Swift code to Dafny and use Hoare logic. As the system we are verifying is complex, our goal is to locate the pieces and target verification in places where formal methods can have the greatest impact; this is a common reality of applying formal methods at scale, see for example (chong2020code, ). Since we are porting the Dafny code to Swift, to reduce the chance of human error, we use a line-by-line port as much as possible and add unit tests for each function to provide confidence that the two implementations are equivalent. We port updated code back to Swift manually because Dafny does not include code generation for Swift.
These efforts paid off: we found and fixed nine bugs in our original implementation; in addition to proving various invariants in the code for our insulin delivery path and security mechanisms. However, the most interesting part was where these bugs were: all of our core algorithms were correct, but we found bugs in seemingly less critical parts of the system, like user settings, which had subtle interactions with the core algorithms.
Our implementation operates on CGM readings, but in our application of formal methods, our goal is to reason about the individuals actual glucose level, taking into account sensor errors. This raises an interesting challenge for formal verification as it requires modeling the true biological state of the user – a parameter that is both unknown at runtime and distinct from measured parameters that are part of the software implementation. We prove that the biological invariant holds, given error bounds set by the CGM manufacturer (garg2022accuracy, ).
To reason about actual glucose levels formally, we use a ghost variable in Dafny. Ghost variables are variables that are only available to the prover, these do not exist in the real software system. By defining the actual glucose levels using a ghost variable, we can include it in preconditions and postconditions to prove properties about our software relative to actual glucose levels.
We believe that having a parameter for the real biological state of the system is an important modeling decision; and it diverges, for example, from previous work that only models the biological data in simulation (zhou2021data, ; zhou2023knowsafe, ). Traditional software verification techniques like model checking (cps_1, ) and deductive verification (cps_3, ) have been explored in the past for hybrid system verification (cps_2, ) that interact with the human body, but they do not completely account for physiological variability between individuals or model inaccuracies during the insulin delivery process. Also, researchers have used real-world invariants for security in the context of autonomous vehicles (choi2018detecting, ), but these were based on quad-copter physics not human physiology and did not include formal verification of their implementation.
@@ 9. Evaluation
In this section, we evaluate GlucOS’s algorithmic and driver security mechanisms. In the next section (Section 10), we evaluate all components of GlucOS in the real world.
@@ 9.1. How effective is GlucOS’s algorithmic security mechanism?
We evaluate GlucOS’s algorithmic security mechanism on 21 virtual humans from an FDA-approved simulator (simglucose, ). GlucOS’s algorithmic security mechanism should protect users against untrusted ML algorithms that administer incorrect insulin doses (overdose or underdose), while minimally impeding algorithms that are efficient at managing T1D. We consider extreme cases of attack in our evaluation. While subtle attacks are also possible, we focus on extreme attacks since they have the most adverse consequences.
Concreteny, we first run simulations with a predictive ML model inspired from prior research (mosquera2023enabling, ), without incorporating our security mechanism to set a baseline of an effctive algorithm in managing T1D. Then, in the same simulation scenarios, we introduce an order of magnitude increase to the insulin doses computed by the baseline algorithm, without incorporating our security mechanism to quantify the risks of untrusted algorithms. We then repeat these experiments with our security mechanism to quantify the impact of the mechanism on correct and incorrect dosing.
The American Diabetes Association (ADA) recommends that individuals with T1D spend at least 70% of their time in the range (TIR) of 70-180 mg/dl to remain healthy (nathan2014diabetes, ; endocrine_tir, ; journal_glycemic_targets, ). Glucose levels less than 70 mg/dl (hypoglycemia) put individuals at risk for acute complications like impaired consciousness, seizure, and death, while levels greater than 180 mg/dl (hyperglycemia) increase risks for long-term vascular damage.

### v4

@@ 4.3. Model principles
The mechanism described above can be implemented with various algorithms. Our contribution is not in the specific algorithms we use, but rather in the insight around insulin delivery being equal for all correct algorithms and how we use this insight for security. In this section, we outline the algorithms we use in our current implementation and define the principles for how to apply security to any predictive insulin dosing algorithms.
@@
Algorithmically, we use a PID controller (franklin2002feedback, ) from feedback control for our reactive safe model, and we use a deep neural network from the literature for meal predictions for our predictive ML model (mosquera2023enabling, ). Although we use this model in our implementation, our system supports any predictive model. The design of the reactive safe model is crucial for algorithmic security, as it serves as the anchor for the security mechanism. The ideal reactive safe model should be safe, easy to understand, and have a formally verified implementation (Section 7). To ensure the safety of this model, we base its calculations solely on facts observed automatically from the CGM and the insulin pump. Given these high-quality sources of data, and a simple insulin absorption physiological model, we can calculate how much insulin an individual needs and how much previously injected insulin has not yet taken effect. From this calculation, we know how much insulin to inject or withhold.
@@
These calculations are standard calculations that people who inject insulin manually use to determine dosing, making them easy for people who manage T1D or their medical care team to understand. The difference in our system is that we run these calculations automatically and every five minutes to adjust to the latest sensor readings. This basic formulation is standard and used in most automated insulin delivery systems.
Our evaluation shows that a benign predictive ML model outperforms the reactive safe model in terms of glucose regulation, both in simulation (Section 9) and for the individual who uses our system for their real-world insulin dosing (Section 10). Our evaluations in simulation also demonstrate the fatal consequences of using an incorrect or malicious ML for insulin dosing, as well as our mechanism’s ability to withstand such models to protect individuals (Section 9).
5. Holistic security: Biological invariant
We recognize that human physiology is complex and it is difficult to ensure that the BioKernel’s reactive safe model, which forms the basis of algorithmic security, can handle all changes to human physiology. Concretely, since the reactive safe model incorporates parameters that represent the user’s physiological state, any divergence in their values could render the model ineffective in protecting the user. With GlucOS, we introduce a biological invariant to detect when such divergence could endanger the user, and shut off automatic insulin dosing to protect them.
From a high level, the biological invariant runs experiments every 5 minutes to empirically measure physiological parameters. These experiments compare the expected drop in glucose from prior insulin doses against the observed change in glucose sensor readings, thereby providing an end-to-end security mechanism. The biological invariant effectively serves as a “catch-all” off-switch for insulin delivery when the BioKernel’s view of the user’s physiology becomes inaccurate. Although our reactive safe model adapts to small changes in physiological parameters by design, the biological invariant detects large changes that operate outside of the bounds of our design.
Insulin sensitivity, which is the amount of glucose absorbed per unit of insulin, is a key physiological parameter that can change. Changes to insulin sensitivity happen from drastic changes to the user’s glucose absorption, such as during exercise, when the body uses glucose without needing insulin (borghouts2000exercise, ), or if the pump insertion site hits a vein and delivers insulin directly into the blood stream (steyn2023accelerated, ). In such cases, the user becomes more sensitive to insulin where the same amount of insulin will facilitate more absorption of glucose than expected, increasing the risk of insulin overdose. For example, if the user becomes twice as sensitive to insulin while jogging, the same insulin dose would decrease their glucose by twice the expected amount, pushing them towards dangerously low glucose (hypoglycemia).
To implement the biological invariant, we compare the actual change in the user’s glucose, against the BioKernel’s calculations of how much their glucose should have changed based on current insulin doses over a fixed time window. If the observed drop in glucose deviates beyond a clinically determined threshold of 30 mg/dl (30mg_1, ; 30mg_2, ) from the calculated drop in glucose within a window of 30 minutes, we consider the biological invariant to have been violated and shut off insulin delivery. Mathematically, we define the biological invariant as follows:
(1) $\Delta G^{C}-\Delta G^{A}\leq T$
Here, $\Delta G^{C}$ denotes the BioKernel’s calculated drop in glucose based on insulin doses over a 30 minute window, $\Delta G^{A}$ denotes the actual drop in the user’s glucose over the same time window, and $T$ is the threshold of 30 mg/dl. Crucially, we note that we cannot precisely calculate $\Delta G^{A}$ or $\Delta G^{C}$ since CGMs and pumps are prone to measurement errors. To account for these errors under standard use, we enforce error bounds on readings from CGMs (garg2022accuracy, ; dexcom_error, ) and records from pumps (Zisser2011OmniPodAccuracy, ) under standard conditions. Based on these bounds, we check the following condition on the BioKernel’s current estimate of insulin sensitivity:
(2) $S\leq\frac{0.9\cdot G_{t+\Delta t}^{M}-1.1\cdot G_{t}^{M}+T}{B\cdot\Delta t-0.99\cdot\Delta IOB^{M}+0.01\cdot\Sigma I_{i}^{\ M}}$
Here, $S$ is the BioKernel’s estimate of the user’s insulin sensitivity, $G_{t}^{M}$ and $G_{t+\Delta t}^{M}$ are readings from the CGM, $\Delta IOB^{M}$ and $\Sigma I_{i}^{M}$ are calculated from the insulin doses recorded by the pump, $\Delta t$ denotes the 30 minute window, and $B$ is the user’s basal rate, a physiological parameter that often remains stable for days and is less prone to change than insulin sensitivity (Lindmeyer2021BasalProfiles, ). We refer to this condition as the implementation invariant and use formal methods to prove that violations to this condition imply violations to the biological invariant (Section 7). Since all quantities in the implementational invariant are observable, we shut off insulin delivery when it is violated. We show how to adopt mathematical models from the literature (simglucose, ; physio2019, ) to derive the implementational invariant in Appendix B.
We resume normal execution if the implementational invariant goes back to being satisfied within two hours, based on clinical guidance for exercise for people living with T1D (ucla_t1d, ; borghouts2000exercise, ). If it remains violated for more than two hours, we alert the user and transition to manual insulin delivery.
Appendix G discusses the less dangerous scenario when the actual change in glucose is higher than the calculated change.
@@ 6. Last line of defense: Humans
In this section, we describe our design and implementation for incorporating humans as a critical component of our multi-layered security system. While our algorithmic security mechanism and biological invariant defenses form the primary barriers against attacks, humans serve as the last line of defense, to take corrective action against hypoglycemia. We discuss our most interesting findings around user preferences and the architectural iterations to our design from real-world experiences.
@@ 6.1. Predictive alerts for humans
When an individual’s glucose levels are not within the range of healthy values, they need to intervene to bring them back in range. Low glucose, or hypoglycemia, occurs from too much insulin, and there is no way to remove insulin that the automated insulin delivery system has already delivered. To recover from hypoglycemia, people need to eat sugar. Given that people with T1D are already active participants in managing diabetes, we focus on how we can use them to create a more comprehensive security system.
@@
We use predictive yet deterministic mobile notifications to alert people about impending hypoglycemia before it happens (Figure 4). We focus on hypoglycemia since it presents the most immediate and dangerous risk. Alerts for hypoglycemia ae not new. All CGM software comes with non-predictive hypoglycemia alerts by default (dexcom_alerts, ) and glucose prediction algorithms have been well covered in the literature (Josep2020, ; Zhang2021, ; arora2023multivariate, ). Our contribution is using predictive alerts before someone reaches hypoglycemia as a security defense.
In a user study (Section 10.4), we reduced the amount of time six people spent in hypoglycemia as a result of using GlucOS, despite them already having CGM alerts set up prior to our study. We highlight that we did not conduct our user study in a controlled environment, but had participants use our app for predictive alerts as they managed their diabetes in the real world. We also highlight that The number of participants in our study is also roughly consistent with recent human factors in computing research (glucomaker, ; moont1d, ; context_aware_t1d, ; t1d_children, ). Our results show that predictive alerts can be effective as a security measure against hypoglycemia.
Using a simple and deterministic algorithm is important since it runs as a part of our trusted computing base. Algorithmically, we predict hypoglycemia using linear regression over the last 20 minutes of CGM readings and extrapolate forward 15 minutes. We trigger alerts if our prediction crossed the hypoglycemia threshold in the next 15 minutes.
The most interesting finding from our study was that personalization was important. We provided people with the ability to set the prediction threshold and specify how often they wanted the alert to repeat while they remained in hypoglycemia. Five of the six participants customized their alerts and the remaining participant noted that the default values we picked coincided with their alerting preferences. In our first implementation we did not provide the ability to customize their settings and we tried to make the usability vs security tradeoff algorithmically and automatically. But in the end, simplifying the system to enable the personalization of alerting policy was important for our participants.
@@ 6.2. Architectural implications
Moving the alerting logic into the BioKernel did have security implications. It adds complexity to our trusted computing base, but we carefully manage this tradeoff. We minimize the added complexity by using traditional, well-understood algorithms for prediction rather than complex ML models and reuse existing functionality within the BioKernel when possible to keep the implementation simple. This approach allowed us to maintain a high level of confidence in the security and correctness of our alerting system while still providing the necessary functionality.
@@ 7. Formal verification
We formally verify the most critical components of GlucOS to prevent unexpected and undesired outcomes (Thomas1994TheSO, ). Specifically, we verify the implementation for: (1) delivering bolus insulin doses (a single dose delivered immediately), (2) delivering basal insulin doses (multiple doses delivered at a constant rate over a specified time period), (3) clamping insulin doses using the reactive safe model, and (4) enforcing the biological invariant. We consider these components to be the most critical since they implement our security mechanisms and directly control insulin delivery.
Formally verifying our GlucOS’s implementation in Swift presented an interesting challenge since automatic translators do not exist between Swift and languages for verification, such as Dafny. While automatic translators exist between Dafny and other languages (such as C) we could not pick those languages over Swift for implementation. This is because Swify provides the most support to build iOS applications and our goal was to build a complete automated insulin delivery application for real-world use.
To overcome this limitation, we manually ported our Swift implementation to Dafny. While manual translation is prone to human error, we took important steps to reduce the likelihood of such errors. Concretely, we ported our Swift code line-by-line to Dafny, preserved variable and function names across implementations (chong2020code, ), and wrote unit tests for each function in both languages to ensure behavioral consistency.
We establish invariants among the aforementioned critical components along the insulin delivery path (Thomas1994TheSO, ). These invariants enforce critical constraints such as maintaining appropriate units insulin doses, disallowing immediately successive bolus doses, ensuring basal doses match doses supported by the pump etc. Collectively these invariants are important to ensure the expected safe operation of the BioKernel. We discover 9 bugs in our original implementation by verifying 31 invariants across 8 functions. Other than the biological invariant (Section 5), some of our other invariants enforced that there should be a gap of at least 4.2 minutes between bolus doses, the amount of insulin on board should also take into account the insulin from basal doses etc. We list all our invariants and bugs in Appendix F.
From our experience with formal verification, we report two key takeaways about our implementation for secure automated insulin delivery. First, we were able to prove the correctness of even those components that involve unobservable quantities (such as the user’s true physiological state in the biological invariant) under standard operation. Second, while we expected most bugs to reside within the implementation of core algorithms, they were fully concentrated among peripheral components such as user settings.
7.1. Capturing true physiological state
Glucose readings from CGMs and insulin doses recorded from pumps do not reflect the user’s true physiological state. This is because CGMs and pumps are subject to measurement errors, making precise glucose concentrations and insulin doses unavailable in Swift.
To model this discrepancy, we use ghost variables in Dafny (Dafny, ) to represent the true glucose values and insulin doses. We then introduce device invariants that bound the maximum measurement error within standard use, i.e., 10% for CGMs (garg2022accuracy, ; dexcom_error, ) and 1% for pumps (Zisser2011OmniPodAccuracy, ). We enforce these bounds as preconditions when verifying the implementation of all our core components. Most importantly, we proved the correctness of the implementational invariant (Equation $(2)$ in Section 5) to verify the biological invariant (Equation $(1)$ in Section5 as a post condition in Dafny).
7.2. Discovering bugs outside the core algorithms
7 out of the 9 bugs we discovered with formal verification were related to issuing insulin delivery commands with invalid parameters. With these bugs, users could specify negative amounts of insulin for basal and bolus doses, set 0-second duration for basal doses, allowing the time to stop a basal dose to be preced the start time of the dose etc. Other bugs involved insfficient error handling, for example when disallowing certain invalid bolus doses. We did not discover any bugs in GlucOS’s core algorithms, such as those used by the reactive safe model. We resolved bugs by updating our Swift implementation to match the formally verified Dafny code.
@@ 9. Evaluation
In this section, we evaluate GlucOS’s algorithmic security mechanism in simulation. In the next section (Section 10), we evaluate all components of GlucOS in the real world. We do not evaluate the biological invariant in simulation since existing simulators do not support changes to human physiology.
@@ 9.1. How effective is GlucOS’s algorithmic security mechanism?
We evaluate GlucOS’s algorithmic security mechanism on 21 virtual humans from an FDA-approved simulator (simglucose, ). The mechanism should protect users against untrusted models that administer incorrect insulin doses (overdose or underdose), while minimally impeding algorithms that are efficient at managing T1D. While subtle attacks are also possible, we focus on extreme attacks in our evaluation since they have the most adverse consequences.
We first run simulations with a predictive ML model inspired from prior research (mosquera2023enabling, ), without incorporating our security mechanism to set a baseline of an effctive algorithm in managing T1D. Then, in the same simulation scenarios, we introduce an order of magnitude increase to the baseline algorithm’s insulin doses to quantify the risks of untrusted algorithms. We then repeat these experiments with our security mechanism to evaluate its impact.
The American Diabetes Association (ADA) recommends that individuals with T1D spend at least 70% of their time in the range (TIR) of 70-180 mg/dl to remain healthy (nathan2014diabetes, ; endocrine_tir, ; journal_glycemic_targets, ). Glucose levels less than 70 mg/dl (hypoglycemia) put individuals at risk for acute complications like impaired consciousness, seizure, and death, while levels greater than 180 mg/dl (hyperglycemia) increase risks for long-term vascular damage. We use TIR as our metric in these evaluations.